OpenStudy (anonymous):

Given the function f(x) = 2(x − 1)2 + 1, indicate the shifts that will affect the location of the vertex and explain what effect they will have. Use complete sentences. f(x+3) f(x) + 3 f(3x) 3•f(x)

OpenStudy (anonymous):

I'm very lost

OpenStudy (anonymous):

your functions is \[f(x)=2(x-1)^{2} +1\] ???

OpenStudy (anonymous):

The function is f(x) = 2(x − 1)^2 + 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

anyone good at math im in 8th anyone above

OpenStudy (danjs):

f(x+3) = 2((x+3) - 1)^2 + 1

OpenStudy (anonymous):

I just want to understand what I'm even doing.

OpenStudy (danjs):

f(x+3) = 2(x+2)^2 +1

OpenStudy (danjs):

see i just put in the (x+3) for the x ?

OpenStudy (anonymous):

yeah

OpenStudy (danjs):

In the form f(x) = a*(x - b)^2 + c the b term shifts the graph sideways the c term shifts the graph up and down the a term streatches the graph or compresses it

OpenStudy (anonymous):

So what tells me how the vertex changes?

OpenStudy (danjs):

So if you put in f(x+3) = 2((x+3) - 1)^2 + 1 you get f(x+3) = 2(x+2)^2 +1

OpenStudy (danjs):

In that case the b term is now -2 (x-(-2))

OpenStudy (danjs):

so you changed the b term by a +3

OpenStudy (anonymous):

alright so what is the vertex in the equation or have we not even found it. I'm terrible at algebra.

OpenStudy (danjs):

it will move the vertex, the whole graph for that matter, sideways to the left by 3 units

OpenStudy (anonymous):

Ohhh okay I get it now.

OpenStudy (danjs):

the standard "Vertex form is" y = a(x-h)^2 +k where (h,k) is the vertex point

OpenStudy (danjs):

so you get the first movement from the f(x+3) ?

OpenStudy (anonymous):

so the question is I have to explain why it affects it and how it affects it.

OpenStudy (danjs):

you are putting in (x+3) into the equation wherever you see x

OpenStudy (danjs):

the standard form has (x - h) so if you add 3, it is the same as (x - (-3)) so that is moving the graph -3 in the x direction

OpenStudy (anonymous):

(The vertex is the minimum or maximum in a quadratic eq) btw

OpenStudy (anonymous):

Okay what tells us how much the vertex changes when we do this

OpenStudy (danjs):

how much the (h,k) changes in y = a(x-h)^2 +k

OpenStudy (anonymous):

okay so now what is the (h,k) in the new equation

OpenStudy (danjs):

We have f(x) = 2(x − 1)2 + 1 f(x+3) = 2((x+3) - 1)^2 +1

OpenStudy (danjs):

so the h term changes from -1 to a +2, so the graph shifts to the LEFT by 3 units

OpenStudy (danjs):

(x - h) = (x -(-2)) = (x+2) the plus 2 means LEFT (x -(-2)) the h term is minus 2

OpenStudy (danjs):

so in summary, f(x+3) shifts the graph to the left 3 units

OpenStudy (anonymous):

Thank you i get it now

OpenStudy (danjs):

the second is easier

OpenStudy (danjs):

f(x) + 3, you are going to add 3 to the k term in y = a(x-h)^2 +k

OpenStudy (danjs):

Shift the graph up by 3 units

OpenStudy (danjs):

f(x) + 3 = 2(x-1)^2 +1 +3 f(x) + 3 = 2(x-1)^2 + 4 (h,k) is the vertex, and the k value increased by 3 units, (SHift UP 3 units)

OpenStudy (danjs):

for f(3x), if you put that into your equation, you can see it will not effect either h or k, the graph will not be shifted, in this case it is compressed

OpenStudy (danjs):

3*f(x) , if you multiply the entire equation by 3, you see it will affect the a term and the K term, the graph will be compressed and shifted upwards

OpenStudy (danjs):

f(x) = a(x-h)^2 +k 3*f(x) = 3a(x-h)^2 +3k